OpenStudy (anonymous):

Ken griffey jr hit 23 homes run in the 50 games of the season if he hit homes run at this space all season how many would he have at the end of the year assume a 168 game schedule?

OpenStudy (anonymous):

(23/50) = (x/168) cross multiply 3864 = 50x x = 77.28 rounded to 77 (sorry about the last one but this one is right)